OpenStudy (anonymous):

What are the solution(s) to the following system of equations? y = x2 + 2 y = x + 2

OpenStudy (anonymous):

Put both equations equal to each other and solve for x.

OpenStudy (anonymous):

\[x ^{2}+2=x+2\] \[x ^{2}-x=0\] x(x-1)=0 x = 0 or x = 1 (0,0) (1,0)

OpenStudy (anonymous):

put the second equation in the 1st or the 1st equation in the 2nd